### v1

# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

Abstract
We present Connection-Aware Motif Sequencing (CamS), a graph-to-sequence representation that enables decoder-only Transformers to learn molecular graphs via standard next-token prediction (NTP). For molecular property prediction, SMILES-based NTP scales well but lacks explicit topology, whereas graph-native masked modeling captures connectivity but risks disrupting the pivotal chemical details (e.g., activity cliffs). CamS bridges this gap by serializing molecular graphs into structure-rich causal sequences. CamS first mines data-driven connection-aware motifs. It then serializes motifs via scaffold-rooted breadth-first search (BFS) to establish a stable core-to-periphery order. Crucially, CamS enables hierarchical modeling by concatenating sequences from fine to coarse motif scales, allowing the model to condition global scaffolds on dense, uncorrupted local structural evidence. We instantiate CamS-LLaMA by pre-training a vanilla LLaMA backbone on CamS sequences. It achieves state-of-the-art performance on MoleculeNet and the activity-cliff benchmark MoleculeACE, outperforming both SMILES-based language models and strong graph baselines. Interpretability analysis confirms that our multi-scale causal serialization effectively drives attention toward cliff-determining differences.
Graph Representation Learning, Autoregressive Modeling, Self-Supervised Learning, Science Foundation Models, Molecular Property Prediction
1 Introduction
Molecular property prediction is a core challenge in drug discovery (wu2018moleculenet) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (khan2025comprehensive) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (wang2024emu3; xia2025nature; zhang2025unigenx). Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (touvron2023llama). However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (xia2025nature).
Conversely, graph-native methods explicitly model topology but typically depart from pure decoder-only NTP, instead adopting hybrid architectures or objectives that involve, to varying degrees, input corruption (e.g., masked node prediction) (li2023knowledge; shehzad2024graph; lu2025uni; kong2025unimomo). While sophisticated strategies—such as incorporating global knowledge nodes or avoiding random masking—have been developed to mitigate the information loss caused by masking (li2023knowledge; rong2020self; liu2024mask; you2020graph; xu2021self), these methods remain fundamentally bound by the corruption-coverage trade-off (wettig2023should). This limitation is particularly acute in activity-cliff scenarios, where a single atomic change triggers a drastic property shift; masking the local neighborhood of a critical functional group effectively removes the precise evidence needed to discern such subtle differences. Standard NTP, by contrast, offers a dense supervision signal without corrupting the visible prefix context (clark2020electra), yet it lacks a representation interface that exposes graph connectivity as effectively as text.
To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design. First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (shibata1999byte; geng2023novo; shen2024graphbpe). This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control. To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism. Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (bundy1984breadth). This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups. Third, and most critically, to construct a hierarchical context, we concatenate subsequences from fine to coarse. This strategy circumvents the trade-off between scales and establishes an Inter-scale Order in the final CamS sequence, enabling high-level motifs to be predicted by conditioning on dense, uncorrupted fine-scale evidence. Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (touvron2023llama). To keep attribution controlled, we inject only a lightweight molecular fingerprint (cereto2015molecular) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven. Across MoleculeNet (wu2018moleculenet) and the activity-cliff stress test MoleculeACE (van2022exposing), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (li2023knowledge). Beyond accuracy, we provide mechanism-level evidence. Targeted ablation studies confirm that multi-scale concatenation is the key driver of performance, while interpretability analysis reveals that the causal serialization effectively drives attention toward cliff-determining structural differences.
Contributions. (1) Methodological Framework: We introduce CamS, a unified graph-to-causal-sequence interface that resolves the conflict between graph topology preservation and scalable autoregressive training, enabling NTP to function as a structure-native objective. (2) Empirical Validation: We demonstrate that CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE benchmarks. Mechanism analysis reveals that the multi-scale causal context explicitly improves attention focus on subtle, activity-determining structural edits. (3) Implementation Recipe: We provide a reproducible pipeline—from the CamS-tokenizer to the CamS-LLaMA—establishing a robust baseline for applying standard foundation model architectures to molecular science.
2 Related Work
2.1 Molecular Property Prediction within Foundation Models Paradigm
Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (awais2025foundation; xia2025nature). One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (xia2025nature; cai2025chemfm). This leverages mature LLM infrastructure and enables straightforward large-scale training. However, SMILES is not structure-native, weakening topology-based signals. String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (xia2025nature). Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (you2021graph; liupre; you2020graph; xu2021self). High-performing methods typically employ encoder-only Graph Transformers (shehzad2024graph) designed for property prediction. While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (li2023knowledge; rong2020self). More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (kong2025unimomo; lu2025uni; zhang2025unigenx).
2.2 Molecular Tokenization Strategies
Within a foundation model paradigm, tokenizer design becomes a key interface for scaling LLMs. In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model. Molecular tokenization spans string-based and graph-based methods (wang2019smiles; liu2023rethinking). String-based schemes operate on SMILES (weininger1988smiles) or related linearizations (krenn2022selfies), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (liu2023rethinking). Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (zhang2021motif); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (li2023knowledge); ring-or-path-based (wollschlager2024expressivity) and tree-based (jin2018junction) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (sun2024representing) or by training a GNN as a tokenizer (liu2023rethinking; luonode). Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (geng2023novo; shen2024graphbpe; kong2022molecule). In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (ji2022relmole; chenhierarchical).
3 Method
Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (touvron2023llama). Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1). Built on CamS, CamS-LLaMA performs large-scale autoregressive pre-training and injects a only lightweight fingerprint prior during downstream fine-tuning (Section 3.2). Finally, we analyze the characteristics and advantages of multi-scale CamS autoregressive modeling by comparing it with Graphformer-style modeling (Section 3.3).
3.1 CamS-Tokenizer
The CamS-tokenizer framework operates in three logical phases to bridge graph topology and scalable NTP. (1) CamS Vocabulary Construction: we maintain a robust vocabulary via BPE-style mining and Single-Atom Vocabulary Closure (SAVC) (Section 3.1.1). (2) Per-Scale Encoding: molecules are partitioned into motif graphs (BPEGraph) and serialized into single-scale CamS subsequences via scaffold-anchored BFS (Section 3.1.2). (3) Cross-Scale Concatenation: subsequences are organized in a fine-to-coarse order to form the final CamS sequence, enabling multiscale autoregressive modeling (Section 3.1.3). Full details are in Appendix A.
3.1.1 Substructure Level: CamS Vocabulary
BPE-style Mining. Following (geng2023novo; shen2024graphbpe), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics. We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs. This yields the BPEMotif set (frequent subgraphs with connectivity patterns) and the base BPE-SAV (Single-Atom Vocabulary with connectivity patterns) naturally captured during mining.
Single-Atom Vocabulary Closure (SAVC). While generation operates within a closed vocabulary, predictive encoding encounters unseen atom states. Standard GraphBPE tokenizers greedily match largest motifs and recursively back off; if a specific atom-connectivity form is absent, it collapses to [UNK]. This erases chemical details (e.g., element type) and creates Causal Information Breakpoints in the autoregressive stream. To prevent this, SAVC constructs the BPE-SAV Back-off via two strategies: (1) enumerating connection-aware tokens for all typical valences, and (2) mapping rare atypical forms of element X to <X_AltForm> rather than <UNK>.
The final CamS Vocabulary $\Sigma$ is the union of BPEMotif, the complete SAV (BPE-SAV $\cup$ Back-off), and special tokens mapped to unique integers. This guarantees that every atom is covered by either a high-level motif or a precise single-atom descriptor (details in Appendix A.1).
3.1.2 Per-Scale Encoding
Given a vocabulary scale $s$ (determined by the iteration count of merge operations), encoding a molecule $M$ involves three stages to generate a single-scale CamS subsequence.
BPEGraph Construction. We apply the prefix of merge operations to $M$, inducing a partition of atoms into non-overlapping fragments. This forms the BPEGraph $G_{s}=(V_{s},E_{s})$, where nodes $V_{s}\subset\Sigma$ are connection-aware motifs/atoms and $E_{s}$ represents chemical bonds between them.
Serialization (Intra-scale Order). To adapt the graph for decoder-only NTP, we serialize $V_{s}$ into an ordered list via Breadth-First Search (BFS). We select the motif with the largest atom count (typically the core scaffold) as the root. This strategy prioritizes global backbone structure before local substitutions (zhang2021motif), establishing a stable Center-to-Periphery causal order.
ID Extraction. Mapping the ordered nodes to their indices in $\Sigma$ yields the single-scale CamS subsequence $X^{(s)}=(x^{(s)}_{1},\dots,x^{(s)}_{L_{s}})$.
3.1.3 Cross-Scale Concatenation
Single-scale tokenization faces a trade-off: coarse scales capture global scaffolds but offer low resolution, while fine scales preserve atomic details but fragment high-level structures (e.g., pharmacophores). CamS resolves this by constructing a multi-scale causal context. We tokenize the molecule at multiple scales (e.g., $s\in\{1\mathrm{K},7\mathrm{K},27\mathrm{K},67\mathrm{K}\}$) by slicing $\mathcal{O}$. The resulting subsequences are concatenated in a fixed Fine-to-Coarse order (Inter-scale Order) to form the final CamS sequence $\mathbf{X}$:
$\mathbf{X}=[$ $\,\texttt{[BOS]},X^{(1\mathrm{K})},\texttt{[CONCAT]},X^{(7\mathrm{K})},\texttt{[CONCAT]},$ (1)
$\,X^{(27\mathrm{K})},\texttt{[CONCAT]},X^{(67\mathrm{K})},\texttt{[EOS]}].$
During training, this CamS token stream allows the model to leverage high-resolution local topology (fine scales) as a prefix to condition the prediction of global scaffolds (coarse scales). This effectively embeds bottom-up structural composition into the autoregressive objective without architectural modifications (analysis in Section 3.3).
3.2 CamS-LLaMA
Figure 1: The overall framework of CamS-LLaMA. (a) CamS-Tokenizer: Molecules are transformed into multi-scale causal sequences. Through Per-scale Encoding, motif graphs are serialized via Scaffold-rooted BFS (Intra-Scale Order) to establish CamS Subsequence, followed by Cross-Scale Concatenation to arrange subsequences into a final CamS Sequence from fine to coarse granularities (Inter-Scale Order). (b) Pre-training Pipeline: The model is pre-trained on CamS Token Lists using a standard LLaMA via NTP, learning to aggregate structural information. (c) Fine-tuning Pipeline: During downstream tasks, the pre-trained weights are transferred, and a lightweight fingerprint prior is injected (via projection, prepending in embedding layer and fusion in classifier or regressor) to take part in the prediction of specific properties.
CamS-LLaMA follows a representation-model co-design principle. Section 3.1 maps a molecular graph $G=(V,E)$ into a multi-scale, connection-aware causal sequence $X=(x_{1},\ldots,x_{T})$ (Eq. (1)) under vocabularies $\Sigma_{1\mathrm{K}}\subset\Sigma_{7\mathrm{K}}\subset\Sigma_{27\mathrm{K}}\subset\Sigma_{67\mathrm{K}}\triangleq\Sigma$, via BFS-ordered motif traversals $\mathcal{T}_{s}(G)$ and token mapping $x^{(s)}_{\ell}=\iota_{s}(\tau^{(s)}_{\ell})\in\Sigma_{s\mathrm{K}}\subseteq\Sigma$. The key point is that this graph-to-causal-sequence interface is intentionally designed to be transparent to a vanilla decoder-only Transformer: it exposes topology through a causal order and a fine-to-coarse hierarchy, so we can train with standard NTP without modifying the LLaMA backbone. Based on this interface, CamS-LLaMA follows the foundation model protocol: (1) self-supervised autoregressive pre-training on CamS sequences (figure 1 bottom right), and (2) supervised fine-tuning with a minimal structural prior (fingerprints) injected only at fine-tuning time (figure 1 bottom left).
3.2.1 Autoregressive Pre-training via NTP
CamS-LLaMA adopts a standard LLaMA-style decoder (touvron2023llama); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification. Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq. (1); we treat each view as one NTP training instance (details in Appendix A.3). Given any view $X=(x_{1},\ldots,x_{T})$, token embeddings form $\mathbf{H}^{(0)}$, which is transformed by $L$ causal self-attention blocks into $\mathbf{H}^{(L)}$; a tied output projection then yields $p_{\theta}(x_{t+1}\mid x_{\leq t})$.
Importantly, CamS changes what the model can condition on under the same NTP objective by reshaping the prefix context through fine-to-coarse concatenation (Eq. (1)). Let $\mathcal{I}_{\mathrm{tok}}\subseteq\{1,\ldots,T-1\}$ be valid prediction positions in $X$ (excluding [BOS], [CONCAT], [EOS]), and $x^{\star}_{t+1}$ be the ground-truth next token. The NTP loss is
$\mathcal{L}_{\mathrm{NTP}}(\theta;X)=-\sum_{t\in\mathcal{I}_{\mathrm{tok}}}\log p_{\theta}(x^{\star}_{t+1}\mid x_{\leq t}).$ (2)
Because $X$ concatenates BFS sequences from fine to coarse scales (Eq. (1)), the conditioning prefix $x_{\leq t}$ has an explicit hierarchy: for $t$ in $X_{67}$ (highest-merge segment), the model conditions on all finer-scale tokens $(X_{1},X_{7},X_{27})$ plus the coarse prefix. Thus, coarse-scale prediction errors propagate through attention to both coarse predecessors and the entire fine-scale prefix, providing a fine-to-coarse causal credit assignment under a fully standard decoder.
On the pre-training corpus $\mathcal{D}_{\mathrm{pre}}$, we minimize
$\min_{\theta}\;\mathbb{E}_{X\sim\mathcal{D}_{\mathrm{pre}}}\big[\mathcal{L}_{\mathrm{NTP}}(\theta;X)\big],$ (3)
learning general-purpose molecular representations purely from CamS sequences, without any downstream labels or extra descriptors.
3.2.2 Fine-tuning with Fingerprint Injection
Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (li2023knowledge). In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
We treat the fingerprint as a virtual global token. For each molecule, a fingerprint vector
$\mathbf{f}\in\mathbb{R}^{D_{\mathrm{fp}}},$ (4)
is projected to the model hidden size:
$\mathbf{f}^{\prime}=W_{\mathrm{fp}}\mathbf{f},\quad W_{\mathrm{fp}}\in\mathbb{R}^{H\times D_{\mathrm{fp}}},$ (5)
and prepended to token embeddings:
$\mathbf{H}^{(0)}_{\mathrm{ft}}=[\,\mathbf{f}^{\prime};\mathbf{e}_{1};\dots;\mathbf{e}_{T}\,]\in\mathbb{R}^{(T+1)\times H}.$ (6)
With a modified causal mask, all positions can attend to this global token. After $L$ layers:
$\mathbf{H}^{(L)}_{\mathrm{ft}}=[\,\mathbf{h}^{(L)}_{0};\mathbf{h}^{(L)}_{1};\dots;\mathbf{h}^{(L)}_{T}\,$]. (7)
We use the global token representation $\mathbf{u}^{(i)}=\mathbf{h}^{(L)}_{0}$ for prediction (a decoder-only analogue of [CLS]).
For classification:
$\begin{gathered}\hat{\mathbf{y}}^{(i)}=\mathrm{softmax}\!\big(W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)})\big),\\ \mathcal{L}_{\mathrm{cls}}(\theta,\phi)=-\frac{1}{B}\sum_{i=1}^{B}\log\hat{\mathbf{y}}^{(i)}_{\,y^{(i)}}.\end{gathered}$ (8)
For regression:
$\begin{gathered}\hat{y}^{(i)}=W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)}),\\ \mathcal{L}_{\mathrm{reg}}(\theta,\phi)=\frac{1}{B}\sum_{i=1}^{B}\big(\hat{y}^{(i)}-y^{(i)}\big)^{2}.\end{gathered}$ (9)
Fine-tuning minimizes
$\min_{\theta,\phi}\;\mathbb{E}_{(X,\mathbf{f},y)\sim\mathcal{D}_{\mathrm{task}}}\big[\mathcal{L}_{\mathrm{cls}/\mathrm{reg}}(\theta,\phi;X,\mathbf{f},y)\big].$ (10)
This keeps pre-training purely sequence/structure-driven, while using fingerprints as a lightweight and controlled prior only when optimizing for downstream targets.
3.3 CamS Representation vs. Graph Transformer
We analyze CamS-LLaMA and graph transformer-style models (e.g., KPGT) within a unified framework of Token-Level Graph-Structured Deep Feature Construction. We show that CamS yields (1) a denser and multi-view direct supervision signal and (2) a hierarchical inductive bias that are mathematically distinct from standard masked-node prediction (MNP). Detailed derivations and supporting discussions are provided in Appendix B.
3.3.1 Unified Framework.
Let $G=(V,E)$ be a molecular graph. Both paradigms construct a token graph $\mathcal{G}_{\mathrm{tok}}=(\mathcal{V}_{\mathrm{tok}},\mathcal{E}_{\mathrm{tok}})$ and perform $L$ layers of propagation:
$\begin{split}\mathbf{H}^{(l+1)}&=\mathrm{Agg}\big(\mathbf{A}^{(l)}\,\mathrm{Val}(\mathbf{H}^{(l)})\big),\\ \mathbf{A}^{(l)}_{ij}&\propto\exp\left(\frac{\mathbf{q}_{i}^{\top}\mathbf{k}_{j}}{\sqrt{d}}+\psi_{ij}\right).\end{split}$ (11)
where $\mathbf{H}^{(l)}$ are token embeddings, $\mathbf{q}_{i},\mathbf{k}_{j}$ are projections of $\mathbf{H}^{(l)}$, and $\psi_{ij}$ encodes the structural bias or causal mask.
Graphformer (Flat/Static Bias). $\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets. $\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (ying2021transformers).
CamS-LLaMA (Hierarchical/Causal Flow). $\mathcal{V}_{\text{tok}}$ comprises multi-scale motifs $\{v^{(s)}\}$, where the same molecule is re-represented across scales. $\psi_{ij}$ is determined by the causal mask $\mathbf{M}_{\text{causal}}$, enforcing a directed flow where coarse-scale tokens attend to fine-scale predecessors. This turns cross-scale aggregation into a learned, anisotropic connectivity along the sequence.
3.3.2 Information-Theoretic Context Analysis.
We contrast NTP with MNP. Let $x_{t}$ be a target token and let $Z_{t}$ denote its uncorrupted evidence set (e.g., the causal history for CamS coarse-scale tokens). MNP observes a stochastically corrupted context $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ via random masking.
Proposition 3.1 (Context Information Inequality).
For any non-trivial masking channel $\mathcal{M}$, we have
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}).$ (12)
Sketch. The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (cover1999elements). In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns. Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (li2023knowledge; rong2020self; liu2024mask). In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss. This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (yang2019xlnet).
3.3.3 Gradient Flow Density.
We define the Direct Supervision Density (SD) as the expected fraction of tokens serving as regression targets per update. For a multi-scale sequence with total length $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$ and mask ratio $\rho$:
$\text{SD}_{\text{NTP}}\approx 1\quad\gg\quad\text{SD}_{\text{MNP}}=\rho.$ (13)
This trade-off (more predicted tokens vs. more corruption) is central to masked modeling (wettig2023should). Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (li2023knowledge). Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (clark2020electra). CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
3.3.4 Hierarchical Inductive Bias.
Graphformers rely on implicit depth and static structural biases to propagate information. In contrast, CamS injects an explicit structural hierarchy via the multi-scale sequence $X=[X_{1},\dots,X_{T_{\text{all}}}]$, and the causal mask ensures coarse-scale predictions have deterministic access to fine-scale details. We hypothesize this is particularly beneficial for activity-cliff regimes, where small local motif edits trigger global property shifts, as it biases the model toward bottom-up composition rules.
4 Experiment
4.1 Experimental Setup
We decouple the pipeline into three stages, using the ChEMBL-34 Dataset for tokenizer training, the Enamine675M Dataset for NTP pre-training, and the MoleculeNet/MoleculeACE Benchmark for downstream fine-tuning and evaluation.
CamS-Tokenizer Training. CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (landrum2021rdkit; ehrlich2011maximum) to extract motifs and attachment patterns, which is difficult to accelerate. We therefore train the tokenizer once on ChEMBL-34 (gaulton2017chembl), a moderately sized yet chemically diverse corpus, at the largest scale (67K). Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements. Smaller scales are obtained without retraining by prefix slicing: each scale uses a prefix $\mathcal{O}_{\leq k}$ together with its induced vocabulary subset (i.e., motifs mined by applying $\mathcal{O}_{\leq k}$), where 1K uses k=0k{=}0, 7K/27K use k=62/210k{=}62/210, and 67K uses k=685k{=}685. Implementation details and full procedures are provided in Appendix A.1.
Pre-training Dataset and Configuration. We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers. For self-supervised pre-training, we use Enamine675M (675M molecules) (shivanyuk2007enamine) and optimize NTP (Section 3.2.1). For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq. (1). We treat each sequence as one training instance and uniformly shuffle all instances, resulting in roughly $5\times$ as many sequences as molecules. All optimization details (schedule, batch size, steps) are summarized in Appendix C.1.
Downstream Benchmarks and Baselines. We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (wu2018moleculenet) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (van2022exposing). To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (li2023knowledge). On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective. On MoleculeACE, we report the benchmark’s comprehensive ML/DL baselines, since activity cliffs emphasize discriminating small structural changes where strong feature-based methods remain competitive. Unless stated otherwise, CamS-LLaMA is kept at $\sim$100M parameters (comparable to KPGT); fingerprints are injected only during fine-tuning as a minimal prior, while pre-training is purely sequence-based.
4.2 Main Results
We summarize the main comparisons in Table 1 (MoleculeNet) and Table 2 (MoleculeACE).
General property prediction (MoleculeNet). Table 1 shows that CamS-LLaMA achieves overall state-of-the-art on MoleculeNet under a comparable model scale to KPGT ($\sim$100M): it slightly improves the average AUROC in classification (0.845 vs. 0.843) and the average RMSE in regression (1.172 vs. 1.175). We also win 6/11 individual tasks (5/8 classification and 1/3 regression), and improve over the previous best self-supervised learning baselines by a clear margin (AVG-AUROC: 0.845 vs. 0.825; AVG-RMSE: 1.172 vs. 1.272). Notably, this is achieved with weaker prior injection: unlike KPGT, which explicitly feeds rich handcrafted descriptors, we keep pre-training purely sequence-based and inject only a lightweight fingerprint prior during downstream fine-tuning. Full per-baseline results and extended tables are provided in Appendix D.
[…]
Unified Takeaway. Taken together, Tables 12 indicate that CamS’s three-stage design (motif mining with single-atom completion $\rightarrow$ causal serialization $\rightarrow$ fine-to-coarse multi-scale concatenation) yields two complementary benefits: (1) general-purpose representation strength, by organizing functional-group semantics, attachment patterns, and topology into an NTP-friendly causal context; and (2) activity-cliff-oriented discriminability, by exposing cliff-driving edits through motif-level abstraction, position-sensitive causal contexts, and redundant multi-scale views. This motivates a targeted interpretability analysis below, where we test whether the model indeed allocates more attention to cliff-differential tokens than to shared scaffolds, and how this preference varies across scale regions.
[…]
Task Previous Best Baselines KPGT CamS-LLaMA
Method Score (Ours)
Classification Tasks (AUROC $\uparrow$)
bace GEM 0.857(0.016) 0.855(0.014) 0.870(0.013)
bbbp GEM 0.895(0.024) 0.908(0.012) 0.942(0.015)
clintox GEM 0.905(0.027) 0.946(0.026) 0.935(0.017)
estrogen GEM 0.894(0.048) 0.906(0.034) 0.917(0.050)
metstab GROV 0.876(0.046) 0.889(0.057) 0.891(0.059)
sider JOAO 0.640(0.012) 0.649(0.011) 0.655(0.016)
toxcast GEM 0.733(0.020) 0.745(0.003) 0.724(0.008)
[…]
AVG GEM 0.825 0.843 0.845
Regression Tasks (RMSE $\downarrow$)
esol GEM 0.803(0.051) 0.804(0.102) 0.761(0.046)
freesolv MolF 2.322(0.613) 2.121(1.025) 2.110(0.959)
lipo GROV 0.625(0.007) 0.600(0.012) 0.645(0.023)
AVG MolF 1.272 1.175 1.172
Table 2: Performance on MoleculeACE (RMSE). AVG: Average RMSE across all 30 tasks. Ranking: Bold (1st), Underline (2nd), Italic (3rd). Metric: $\downarrow$ denotes lower is better. Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning). Reference & Abbreviations: GMAE: GraphMAE (hou2022graphmae); MBRT: MoleBERT (xia2023mole); GROV: GROVER (rong2020self); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (van2022exposing). Descriptors: E: ECFP (rogers2010extended); M: MACCS (durant2002reoptimization).
Task Previous Best Baselines KPGT CamS-LLaMA
Method Type Score (Ours)
Regression Tasks (RMSE $\downarrow$)
[…]
4.3 Interpretability: Attention on Activity Cliffs
Interpretability setup and metric. To explain the activity-cliff advantage in Table 2, we test whether CamS-LLaMA preferentially attends to cliff-driving structural differences. Following the MoleculeACE protocol, we construct activity-cliff pairs on the test sets and identify differential vs. shared atom-level activity-cliff point; we then map them to tokens within each scale region of the concatenated CamS sequence (Appendix E). Given the final-layer attention distribution, we compute mean attention on differential tokens (MDTA) and shared tokens (MSTA) within a region and report the relative preference $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100$, evaluated for each scale region and for the full concatenated sequence. We report the full attention metrics (MDTA/MSTA/Rel-DTAP) in Appendix D.
Findings and implications. Table 3 shows three patterns. (1) Overall preference is positive: on the concatenated sequence, Rel-DTAP is consistently >0>0 (14.79% w/o fingerprint; 14.23% w/ fingerprint), indicating that the model, on average, prioritizes cliff-driving differences over shared scaffolds. (2) Preference is scale-dependent: coarse regions (27K/67K) exhibit much stronger preference (24.33%/25.87% w/o fingerprint) than the fewest-merge region (1K, 8.08%), suggesting that coarse motifs provide a more salient interface for highlighting differential substructures, while fine scales contribute weaker but complementary local evidence. (3) Fingerprint acts as a stabilizer rather than a uniform booster: although fingerprint slightly reduces the overall preference, it corrects the anomalous 7K behavior (from -0.18% to 2.10%), consistent with fingerprint providing global guidance that resolves ambiguity at mid-scale tokenizations. Together, these results support the cliff-specific advantage of CamS: multi-scale causal serialization does not merely improve accuracy, but also steers attention toward structurally meaningful differences that drive activity cliffs.
Table 3: Relative Differential-Token Attention Preference (Rel-DTAP). Within each scale region of the concatenated multi-scale sequence, we compute per-molecule $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100\%$, then average over all molecules from activity-cliff pairs (anchor and partner counted separately). Positive values mean higher per-token attention on differential tokens than on shared tokens.
Scale region Without FP With FP
1K 8.08% 4.72%
7K -0.18% 2.10%
27K 24.33% 23.60%
67K 25.87% 24.00%
CONCAT (all scales) 14.79% 14.23%
4.4 Ablation Study
Ablation setup. To isolate the contribution of each core component, we evaluate three variants against the full model (Table 4): (1) w/o FP removes fingerprint injection during fine-tuning; (2) Vocab-1K Only and (3) Vocab-67K Only remove multi-scale concatenation by restricting both pre-training and fine-tuning to a single scale. For fair comparison, single-scale variants keep the pre-training budget matched to the full model (details in Appendix C.1).
Multi-scale concatenation is indispensable. The full model consistently outperforms both single-scale variants across all benchmarks (e.g., MolNet-Cls 0.845 vs. 0.833/0.818; MolNet-Reg 1.172 vs. 1.215/1.329; MolACE 0.624 vs. 0.641/0.649), confirming that multi-scale concatenation is not a minor engineering choice but the key mechanism that lets coarse-scale reasoning build on fine-scale evidence within one causal context. This directly supports our central claim: multi-scale causal serialization is the representation-level ingredient that makes graph structure learnable by a vanilla decoder-only NTP backbone.
Why 1K-only can beat 67K-only under a fixed NTP budget. Under single-scale training, Vocab-67K Only is consistently weaker than Vocab-1K Only, especially on regression (1.329 vs. 1.215 RMSE on MolNet-Reg). We interpret this as an over-compression effect: at 67K, sequences for small molecules become very short and each token packs high compositional complexity, which reduces the number of NTP supervision steps per molecule and makes causal dependency learning harder under limited model capacity and training steps. In contrast, 1K produces longer sequences that provide denser autoregressive supervision and more stable causal credit assignment, albeit at the cost of fragmenting scaffold-level semantics. Concat resolves this trade-off by retaining the dense supervision of fine scales while reintroducing scaffold-level abstraction at coarse scales, without committing to a single “best” vocabulary size.
[…]
Table 4: Ablation study and ranking comparison. Scores represent the average performance across all tasks in the respective benchmarks from Tables 1 and 2. Rank: Global ranking of the specific variant when pooled with the complete set of baseline models. Previous Best: The highest-performing baseline excluding KPGT. Abbreviations: MolF: MolFormer
Method MolNet-Cls ($\uparrow$) MolNet-Reg ($\downarrow$) MolACE ($\downarrow$)
Score Rank RMSE Rank RMSE Rank
CamS-LLaMA (Full) 0.845 1 1.172 1 0.624 1
– w/o FP 0.838 3 1.195 3 0.650 5
– Vocab-1K Only 0.833 4 1.215 4 0.641 3
– Vocab-67K Only 0.818 6 1.329 7 0.649 4
KPGT 0.843 2 1.175 2 0.633 2
[…]
5 Conclusion
We introduced CamS, a graph-to-sequence interface that enables standard decoder-only Transformers to learn molecular topology via Next Token Prediction. By unifying motif mining, causal serialization, and fine-to-coarse concatenation, CamS constructs a hierarchical context where global scaffolds are conditioned on dense local evidence. The resulting CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE, surpassing strong graph baselines while relying on minimal priors. Mechanism analysis confirms that this multi-scale causal design is crucial, as it explicitly drives attention toward cliff-determining structural differences. Our work validates that with the right representation interface, generic autoregressive foundation models can serve as powerful, structure-native engines for molecular science.
Appendix A CamS-tokenizer and Graph-to-Sequence Construction
Overview. This section corresponds to the CamS-tokenizer description in Sec. 3.1 (including Secs. 3.1.1, 3.1.2, and 3.1.3) and the multi-view training-instance construction in Sec. 3.2.1. It supplements the main text with implementation-level details of (1) GraphBPE merge learning and Single-Atom Vocabulary Closure (SAVC)(Algs. 1 and Algs. 2), (2) connection-aware motif vocabulary construction and encoding-time unknown recovery (Algs.3 and Algs.4), (3) deterministic scaffold-rooted BFS serialization (Algs. 5), and (4) fine-to-coarse multi-scale concatenation together with NTP loss masking and length handling (Algs. 6)
A.1 Tokenizer Vocabulary Mining and Single-Atom Coverage (GraphBPE + SAVC)
Notation and objects.
We start from an RDKit molecule and construct the atom graph $G_{0}=(V_{0},E_{0})$, where each node is an atom and each edge is a chemical bond. A GraphBPE merge list is an ordered sequence of operations $\mathcal{O}=(o_{1},\dots,o_{K})$, where each operation $o_{t}$ is a canonical fragment code extracted from the union of two adjacent nodes (details below). Given a prefix $\mathcal{O}_{\leq k}$, applying merges on $G_{0}$ yields a motif graph $G=(V,E)$, where each motif node $v\in V$ stores the covered atom set $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
Connection-aware motif representation.
Each motif token is stored in a pair form $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$. The first records the fragment SMILES of the core atoms (no connectors), while the second inserts wildcard connector atoms * at all attachment sites together with explicit bond types, making the token identity connection-aware. We materialize these motif pairs by replaying $\mathcal{O}_{\leq k}$ on a tokenizer corpus and inserting connector * atoms for every cut bond (Alg. 3).
Single-atom vocabulary closure (SAVC).
A key issue for property prediction is encoding coverage: if a rare but meaningful “single-atom + attachment-pattern” form is absent from the mined vocabulary, it would collapse to [UNK] and erase fine-grained distinctions. SAVC addresses this by (1) enumerating common connection-aware single-atom motifs for typical valence states, and (2) adding an element-wise fallback token X_AltForm to represent rare/atypical forms. At encoding time, if a queried motif contains exactly one core atom of element $X$ but the exact form is missing, we back off to X_AltForm instead of [UNK] (Alg. 2).
Encoding-time unknown recovery.
Encoding greedily applies $\mathcal{O}_{\leq k}$ and then maps each resulting motif to its token id. If a motif is still unknown after canonicalization, we recursively split it along its stored merge tree until known motifs are obtained; the recursion stops at single-atom leaves where SAVC backoff applies (Alg. 4).
Algorithm-to-component map.
Alg. 1 learns the ordered merge list $\mathcal{O}$ (Sec. 3.1.1). Alg. 3 materializes the connection-aware motif vocabulary by replaying $\mathcal{O}_{\leq k}$ and inserting connector atoms *. Alg. 2 specifies SAVC enumeration and X_AltForm backoff. Alg. 4 describes encoding-time motif extraction and recursive split for unknown recovery.
A.2 Graph-to-Causal-Sequence Serialization (Traversal and Root Selection)
Motif graph.
After applying $\mathcal{O}_{\leq k}$ to $G_{0}$, we obtain a motif graph $G=(V,E)$ where each node $v\in V$ corresponds to a merged fragment with $\mathrm{atom\_indices}(v)\subseteq V_{0}$. An edge $(u,v)\in E$ indicates that at least one original bond in $G_{0}$ connects atoms across the two fragments.
Scaffold-rooted BFS order.
CamS serializes $G$ into a token sequence by a deterministic scaffold-rooted BFS: we select as root the motif with the largest number of covered atoms (ties broken deterministically, e.g., by smaller node id), and traverse the motif graph breadth-first with a fixed neighbor ordering (e.g., ascending node id). The visitation order defines the final token order, implemented by relabeling motif nodes so that ids $0,1,\dots$ encode the BFS order (Alg. 5). This yields a stable core-to-periphery causal order aligned with prefix-based NTP (Sec. 3.1.2).
Why no explicit edges in the sequence.
We do not append edge/direction tokens: each motif token already includes explicit * attachment sites and bond types, so the attachment pattern is part of the token identity (Alg. 3). The BFS order then provides the autoregressive causal structure at the sequence level.
A.3 Multi-scale Concat Construction and Training-Instance Generation
Prefix slicing and scale definition.
GraphBPE yields a single merge list $\mathcal{O}=(o_{1},\dots,o_{K})$. Each vocabulary scale corresponds to a prefix $\mathcal{O}_{\leq k_{s}}$, which induces a sub-vocabulary $\Sigma_{\leq k_{s}}$ and a tokenizer $\iota_{s}$. This “train once, slice many” construction reuses the same learned merge statistics across all scales (Sec. 3.1.3).
Per-molecule views and fine-to-coarse concatenation.
For a molecule $s$, we generate $S$ single-scale views $\{X_{s_{1}},\dots,X_{s_{S}}\}$ by encoding under tokenizers ordered from fine to coarse. We then build one concatenated view by inserting [CONCAT] as separators and keeping only one global [BOS] and [EOS]. Alg. 6 provides the pseudo-code used in implementation (Sec. 3.2.1).
Loss masking and length handling.
For NTP training, we exclude [BOS], [EOS], and [CONCAT] positions from prediction targets (typically by setting labels to -100 at these indices). Sequences are padded with [PAD] up to the maximum length; if truncation is needed, we keep boundary tokens [BOS] and [EOS] and truncate the interior span consistently across views.
Algorithm 1 GraphBPE merging-operation learning (ordered merge list $\mathcal{O}$)
1:Input: tokenizer corpus $\mathcal{D}$ (RDKit molecules), max iterations $K$, min frequency $f_{\min}$
2:Output: ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K^{\prime}})$
3: Initialize each molecule $m\in\mathcal{D}$ with a working graph $G_{m}$ where every atom is a node and $\mathrm{atom\_indices}(i)=\{i\}$
4: Compute initial statistics: for each $m$ and each edge $(u,v)\in E(G_{m})$, set $c\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(u)\cup\mathrm{atom\_indices}(v))$ and increment $\mathrm{stats}[c]$
5:for $t=1$ to $K$ do
6:  Select $o_{t}\leftarrow\arg\max_{c}\;(\mathrm{stats}[c],\,c)$ (tie-break by lexicographic $c$ as in code)
7:  if $\mathrm{stats}[o_{t}]<f_{\min}$ then
8:   break
9:  end if
10:  Append $o_{t}$ to $\mathcal{O}$
11:  Apply $o_{t}$ to all molecules that contain it: in each affected $G_{m}$, merge every edge $(u,v)$ whose union-fragment code equals $o_{t}$
12:  Update local pair statistics around merged nodes (remove old incident pairs; add new pairs)
13:  Set $\mathrm{stats}[o_{t}]\leftarrow 0$ (avoid re-selecting the same operation)
14:end for
Algorithm 2 Single-Atom Vocabulary Closure (SAVC) with AltForm backoff (implementation-level)
[…]
2:Output: basic single-atom token list $\Sigma_{\mathrm{atom}}$
3:for each element $X\in\mathcal{E}$ do
4:  Add X (standalone single atom) and X_AltForm (rare/atypical fallback)
5:  for each $v\in\mathrm{Val}(X)$ do
6:   for each attachment degree $d\in\{1,\dots,v\}$ do
7:    for each bond-type tuple $(b_{1},\dots,b_{d})\in\mathcal{B}^{d}$ do
8:     Construct a single-atom fragment SMILES with $d$ wildcard substituents (use branching if $d>1$), e.g. *X(*)(*)* with explicit bonds
9:     Add this motif_with_conn into $\Sigma_{\mathrm{atom}}$
10:    end for
11:   end for
12:  end for
13:end for
14:Encoding-time backoff (MotifVocab.__getitem__): if a queried motif SMILES is not found after RDKit canonicalization, and it contains exactly one core atom (non-H, non-*), return the id of X_AltForm; otherwise return [UNK] (code uses id =1=-1 as unknown sentinel).
Algorithm 3 Materializing the connection-aware motif vocabulary (motif mining + SAVC integration)
1:Input: tokenizer corpus $\mathcal{D}$ (SMILES), operation prefix $\mathcal{O}_{\leq k}$, special tokens $\{\texttt{[PAD]},\texttt{[BOS]},\texttt{[EOS]},\texttt{[UNK]}\}$
2:Output: motif-pair list $\mathcal{V}=\{(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\}$ and token set $\Sigma$
3:for each SMILES $s\in\mathcal{D}$ do
4:  Parse $s$ to molecule $m$; build $G_{0}=(V_{0},E_{0})$ with atom/bond features
5:  Apply merges in order $\mathcal{O}_{\leq k}$ on a copy of $G_{0}$ to obtain motif graph $G=(V,E)$ where each node $v$ stores $\mathrm{atom\_indices}(v)\subseteq V_{0}$
6:  Assign each atom $i\in V_{0}$ to its motif node $\mathrm{bpe\_node}(i)\in V$
7:  for each original bond $(i,j)\in E_{0}$ do
8:   if $\mathrm{bpe\_node}(i)\neq\mathrm{bpe\_node}(j)$ then
9:    Add two connector nodes $i^{\star},j^{\star}$ with smarts=*; attach $i^{\star}$ to $i$ and $j^{\star}$ to $j$
10:    Set the bond types of $(i,i^{\star})$ and $(j,j^{\star})$ equal to the original bond type of $(i,j)$
11:    Include $i^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(i))$ and $j^{\star}$ in $\mathrm{atom\_indices}(\mathrm{bpe\_node}(j))$
12:   end if
13:  end for
14:  for each motif node $v\in V$ do
15:   $v_{\mathrm{noConn}}\leftarrow\mathrm{FragSmiles}(m,\mathrm{atom\_indices}(v)\cap V_{0})$
16:   $v_{\mathrm{withConn}}\leftarrow\mathrm{GraphSmiles}(G_{0}[\mathrm{atom\_indices}(v)])$ (contains * connectors)
17:   Add pair $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$ into a frequency counter
18:  end for
19:end for
20: Prepend special tokens; then integrate the SAVC basic vocabulary (Alg. 2) into the pair list
Algorithm 4 GraphBPE encoding with unknown-motif recovery (recursive split)
1:Input: SMILES $s$, operations $\mathcal{O}_{\leq k}$, motif vocabulary $\Sigma$ (lookup by canonical SMILES)
2:Output: ordered motif list $\big[(v_{\mathrm{noConn}},v_{\mathrm{withConn}})\big]$
3:if $s$ contains “.then
4:  Split $s$ into fragments and encode each fragment independently; concatenate motif lists
5:end if
6: Build initial atom graph $G_{0}$ and initialize a binary merge tree at each atom node
7: Apply $\mathcal{O}_{\leq k}$ in order: for each code $o$, merge every edge whose union-fragment code equals $o$ (merge also updates the node’s merge tree)
8: Relabel nodes to integers and reorder them by scaffold-rooted BFS (Alg. 5)
9: Add connector * nodes for every cut bond to build connection-aware subgraphs (Alg. 3)
10: Collect motif tokens in BFS order; if a motif is not in $\Sigma$, mark it as unknown
11:while unknown motifs exist do
12:  Pop one unknown motif node and recursively split it into its two children in the stored merge tree
13:  For each child: if it is in $\Sigma$, emit it; else split further; stop at leaves (single-atom motifs), where SAVC backoff applies
14:end while
Algorithm 5 Scaffold-rooted BFS serialization (root selection + neighbor order)
1:Input: motif graph $G=(V,E)$ with node attribute $\mathrm{atom\_indices}(v)$
2:Output: a deterministic BFS ordering $\pi$ (used to relabel nodes)
3: Select root $r\leftarrow\arg\max_{v\in V}\big(|\mathrm{atom\_indices}(v)|,\;-\mathrm{id}(v)\big)$ (tie-break: smaller id)
4: Initialize queue with $r$, mark $r$ visited
5:while queue not empty do
6:  Pop front node $u$, append $u$ to ordering $\pi$
7:  Let $\mathcal{N}(u)$ be neighbors of $u$ not yet visited, sorted by ascending node id
8:  Push neighbors in $\mathcal{N}(u)$ into the queue and mark them visited
9:end while
10: Relabel nodes by $\pi$ so that the first visited node becomes id 0, the second becomes id 11, etc.
Algorithm 6 Multi-scale CamS concatenation and training-view generation
[…]
2:Output: single-scale views $\{X_{s_{j}}\}_{j=1}^{S}$ and concatenated view $X_{\mathrm{concat}}$
3:for $j=1$ to $S$ do
4:  $X_{s_{j}}\leftarrow\iota_{s_{j}}.\mathrm{encode}(s).\mathrm{ids}$ (already includes [BOS] and [EOS])
5:end for
6: Build concatenated ids:
7:  $X_{\mathrm{concat}}\leftarrow[\texttt{bos\_id}]$
8:for $j=1$ to $S$ do
9:   Append the inner ids of $X_{s_{j}}$ (drop its BOS/EOS): $X_{s_{j}}[2{:}{-}1]$
10:  if $j<S$ then
11:    Append concat_id
12:  end if
13:end for
14:  Append eos_id
15:NTP loss masking: exclude [BOS], [EOS], [CONCAT] positions from prediction targets (implementation typically sets labels to -100 on these ids).
16:Padding/truncation: pad with [PAD] to max length; if truncating, keep the boundary tokens [BOS] and [EOS].
Appendix B Theoretical Derivations
Overview. This section corresponds to the theoretical analysis in Sec. 3.3 and the supporting arguments referenced in Secs. 3.1 and 3.2.1. It supplements the main text with derivation-level details of (1) the DPI-based proof for the information loss induced by stochastic corruption (Proposition 3.1; Sec. B.1), (2) the direct supervision density / target-coverage counting under NTP vs. masked modeling in the multi-scale setting (Sec. B.2), and (3) the explicit structural-bias formulation that contrasts graph-transformer static bias with CamS causal constraint plus learned aggregation (Eq. (11); Sec. B.3).
B.1 Proof of Proposition 3.1 (Context Information)
We formalize the conditioning-information gap between (1) uncorrupted evidence and (2) randomly masked evidence.
Setup.
For predicting token $x_{t}$, let $Z_{t}$ denote an unmasked evidence set (a collection of tokens) and let $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ be its stochastically masked version under a masking channel $\mathcal{M}$ (e.g., random node/substructure masking used in MLM/MNP). This yields a Markov chain
$x_{t}\;\longleftrightarrow\;Z_{t}\;\longrightarrow\;\tilde{Z}_{t}.$ (14)
Proof.
[…]
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}),$ (15)
which proves Proposition 3.1. ∎
Remark (MLM limitations in NLP).
XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (yang2019xlnet). Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (clark2020electra).
Why this matters more on graphs (evidence-pattern uncertainty).
On graphs, the most predictive evidence for a token often lies in its local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings). Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance. This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (li2023knowledge), (2) masking local subgraphs and predicting contextual properties (GROVER) (rong2020self), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (liu2024mask). These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (wettig2023should).
CamS implication.
For CamS, coarse-scale token prediction under NTP conditions on an uncorrupted fine-to-coarse history by construction. Thus, relative to MNP which observes $\tilde{Z}_{t}$, CamS provides a higher-SNR conditioning set $Z_{t}$ that avoids stochastic evidence deletion. This does not claim that an arbitrary prefix always dominates the full bidirectional context; rather, it claims that masking as corruption strictly reduces the information of whatever evidence set is available, and CamS is designed so that the evidence set for coarse tokens is both uncorrupted and structurally aligned with the fine-grained neighborhood evidence.
B.2 Direct Supervision Density Analysis (Multi-Scale)
As a heuristic proxy for the richness of the learning signal, we compare the density of direct token-level supervision signals and how they accumulate over optimization.
Multi-scale notation.
Let the multi-scale CamS sequence consist of $S$ scales with lengths $\{T_{s}\}_{s=1}^{S}$ and total tokens $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$.
Where gradients originate.
Let $\mathcal{L}$ be the token-level loss.
  • MNP: Direct loss terms (hence gradient sources) occur only on masked indices $i\in\mathcal{M}$; unmasked tokens receive gradients only indirectly via attention coupling.
  • NTP: Direct loss terms occur at (almost) all positions $t\in\{1,\dots,T_{\text{all}}-1\}$.
Direct supervision density and accumulation.
We define the Direct Supervision Density (SD) as the expected fraction of tokens that serve as prediction targets per update:
$\text{SD}_{\text{NTP}}\approx 1,\qquad\text{SD}_{\text{MNP}}=\rho.$ (16)
Equivalently, over $K$ optimization steps, the number of direct token-level targets scales as
$\mathcal{R}_{\text{NTP}}\propto K\cdot T_{\text{all}},\qquad\mathcal{R}_{\text{MNP}}\propto K\cdot\rho\,T_{\text{all}}.$ (17)
Thus, NTP provides about $1/\rho$ times more direct targets than MNP (e.g., $6.7\times$ when $\rho=0.15$).
Trade-off and practical masking rates.
Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (wettig2023should). In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (li2023knowledge). More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (rong2020self).
Why dense (and repeated) supervision can help.
Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (clark2020electra). CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales). This reduces sensitivity to any single tokenization granularity and makes the effective supervision more robust.
B.3 Structural Bias Formulation
We distinguish the structural injection mechanism in Eq. (11).
Graphformer (Hard Static Bias).
The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (ying2021transformers):
$\psi_{ij}^{\text{Graph}}=b_{\phi(d_{G}(i,j))}.$ (18)
CamS (Hard Causal Constraint + Soft Learned Aggregation).
CamS uses a hard causal mask to enforce fine-to-coarse information flow:
$\psi_{ij}^{\text{CamS}}=\begin{cases}0&\text{if }(i\!\rightarrow\!j)\text{ is allowed by }\mathbf{M}_{\text{causal}}\\ -\infty&\text{otherwise.}\end{cases}$ (19)
Within the permitted region, connectivity is content-adaptive via $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$; hence coarse-scale tokens learn which fine-scale neighborhood details to aggregate, rather than uniformly mixing neighbors at a given distance.
Appendix C Training and Implementation Details
Overview. This section corresponds to the training and evaluation protocols in Secs. 3.2.1 and 4. It supplements the main text with implementation-level details of (1) the shared training framework and parallelization setup for pre-training vs. fine-tuning (Sec. C.1), (2) the complete pre-training configuration on Enamine675M (MultiScaler3375M) (Table 5), (3) the fixed fine-tuning setup used across MoleculeNet and MoleculeACE (Table 6), and (4) the unified hyperparameter search space (including strategy-specific grids) (Table 7).
C.1 Pre-training and Fine-tuning Hyperparameters
Implementation and framework.
All experiments are implemented in PyTorch using HuggingFace Transformers. Pre-training uses the built-in Trainer with DeepSpeed ZeRO-2 on multi-GPU data parallelism; fine-tuning is run as single-GPU jobs (one job per GPU) using the same Trainer API (or strategy-specific subclasses).
Pre-training on Enamine675M (MultiScaler3375M).
We pre-train a 16-layer LLaMA-style decoder initialized from a local config (i.e., random initialization; no pretrained weights are loaded). The training data are pre-tokenized and stored as HuggingFace datasets on disk with train/validation splits. We use FP16 training with a cosine learning-rate schedule and warmup, and evaluate/save checkpoints periodically during training.
Fine-tuning (MoleculeNet and MoleculeACE).
We fine-tune the pretrained backbone with fingerprint injection (fp_dim=2048) and a lightweight supervised head. All downstream runs use FP16, evaluate every epoch, and select the best checkpoint by the primary validation metric (higher-is-better for AUROC; lower-is-better for RMSE). We also tested two dropout application schemes in the MoleculeACE fine-tuning pipeline: (1) head-only dropout (applied only in the supervised head), and (2) backbone-synced dropout (also overriding backbone attention_dropout/hidden_dropout and all nn.Dropout modules). Both schemes share the same dropout-rate grid in Table 7.
Ablation compute budget.
For ablations (e.g., single-scale tokenization or removing fingerprint injection), we keep the optimizer, LR schedule, batch size, and training length the same as the corresponding main setting, and only change the ablated component.
Table 5: Pre-training hyperparameters on Enamine675M (MultiScaler3375M).
Setting Value
Backbone config 16 layers; hidden size 720; 8 heads; MLP 2880
Context length 4096 (max_position_embeddings)
Initialization new model from local config (no pretrained weights)
Tokenizer CamS-tokenizer; embedding resized to tokenizer.vocab_size
Training instances Multi-scale tokenized corpus (MultiScaler3375M; 5 views per molecule)
Parallelism torchrun 8 processes + DeepSpeed ZeRO stage 2
Precision FP16 (fp16=true in Trainer and DeepSpeed)
Micro-batch size 256 per GPU
Gradient accumulation 2
Global batch size 4096 (256 $\times$ 8 $\times$ 2)
Optimizer AdamW (Transformers Trainer default)
Learning rate $1\times 10^{-4}$
Weight decay $1\times 10^{-2}$
LR schedule cosine; warmup_steps =5000=5000
Gradient clipping 1.0 (DeepSpeed gradient_clipping)
Training length 10 epochs in planning (early stop after 1/10 epoch in actrual excution)
Eval / save / log eval_steps =33000=33000; save_steps =33000=33000; logging_steps =3300=3300
Checkpoint cap save_total_limit =100=100
Table 6: Fixed fine-tuning setup for MoleculeNet and MoleculeACE.
Setting MoleculeNet MoleculeACE
Task type classification / regression (per dataset) regression
Data split provided scaffold splits (same as KPGT), fold $\in\{0,1,2\}$ provided scaffold splits (same as KPGT)
Epochs 80 (search phase) 50
Batch size 32 per GPU (default) 32 per GPU (default; a small subset rerun with 64)
Grad accumulation 1 (default) 1(default)
Optimizer AdamW (Trainer default) AdamW (Trainer default)
LR schedule constant_with_warmup, warmup_ratio =0.1=0.1 constant_with_warmup, warmup_ratio =0.1=0.1
Evaluation per epoch per epoch
Primary metric AUROC / macro-AUROC; RMSE for regression tasks RMSE (overall_RMSE)
Table 7: Fine-tuning hyperparameter search space (shared).
Hyperparameter Candidates
Base grid (48 combos) learning_rate $\in\{10^{-6},3\!\times\!10^{-6},10^{-5},3\!\times\!10^{-5}\}$
dropout_rate $\in\{0,0.05,0.1,0.2\}$; weight_decay $\in\{0,10^{-6},10^{-4}\}$
Fine-tune strategy normal, l2sp, llrd, flag, reinit
LLRD layer_decay $\in\{0.95,0.9,0.85,0.8\}$
L2SP weight $\in\{10^{-1},10^{-2},10^{-3},10^{-4}\}$
FLAG $\epsilon\in\{10^{-3},3\!\times\!10^{-3},5\!\times\!10^{-3},10^{-2}\}$
Reinit reinit_num_layers $\in\{1,2,3,4\}$
Appendix D Detailed Experimental Results
Overview. This section corresponds to the benchmark results summarized in Sec. 4. It supplements the main text with the full detailed tables of (1) MoleculeNet overall performance of all baseline and their rank comparison (extended version of Table 1; Table 8), and (2) MoleculeACE results overall performance of all baseline and their rank comparison (extended version of Table 2; Table 9).
Table 8: Full Performance comparison on MoleculeNet benchmark. (Extended version of Table 1). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Baseline Reference: Infomax (velivckovic2018deep); Edgepred (hamilton2017inductive); Attribute Masking (Masking) (hu2020strategies); Context Prediction (Contextpred) (hu2020strategies); GraphLoG (xu2021self); GraphCL (you2020graph); JOAO (you2021graph); GROVER (rong2020self); 3DInfomax (stark20223d); GraphMVP (liu2016pyridazinone); ImageMol (zeng2022accurate); MolFormer (wu2023molformer); GEM (fang2022geometry); GraphMAE (hou2022graphmae); MoleBert (xia2023mole); KPGT (li2023knowledge)
Classification Tasks Regression Tasks
Method AVG-AUROC $\uparrow$ Rank Method AVG-RMSE $\downarrow$ Rank
CamS-LLaMA 0.845 1 CamS-LLaMA 1.172 1
KPGT 0.843 2 KPGT 1.175 2
CamS-LLaMA (w/o FP) 0.838 3 CamS-LLaMA (w/o FP) 1.195 3
CamS-LLaMA (Vocab-1K Only) 0.833 4 CamS-LLaMA (Vocab-1K Only) 1.215 4
GEM 0.825 5 MolFormer 1.272 5
CamS-LLaMA (Vocab-67K Only) 0.818 6 GEM 1.285 6
GROVER 0.818 7 CamS-LLaMA (Vocab-67K Only) 1.329 7
ContextpredSup 0.807 8 GROVER 1.332 8
3DInfomax 0.804 9 3DInfomax 1.400 9
ImageMol 0.802 10 ContextpredSup 1.414 10
MoleBERT 0.802 11 ImageMol 1.501 11
MaskingSup 0.799 12 MoleBERT 1.559 12
EdgepredSup 0.795 13 Contextpred 1.563 13
InfomaxSup 0.794 14 EdgepredSup 1.576 14
JOAO 0.790 15 MaskingSup 1.620 15
Contextpred 0.790 16 GraphMVP 1.647 16
GraphMAE 0.788 17 InfomaxSup 1.661 17
Edgepred 0.785 18 GraphLoG 1.663 18
GraphCL 0.774 19 Edgepred 1.674 19
Infomax 0.773 20 Masking 1.697 20
Masking 0.773 21 GraphMAE 1.716 21
GraphLoG 0.769 22 Infomax 1.770 22
GraphMVP 0.769 23 GraphCL 1.822 23
MolFormer 0.744 24 JOAO 1.960 24
Table 9: Full Detailed results on 30 MoleculeACE activity-cliff tasks (Regression, RMSE $\downarrow$). (Extended version of Table 2). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Descriptors: ECFP (rogers2010extended); MACCS (durant2002reoptimization); PHYSCHEM (walters2002prediction); WHIM (kubinyi19933d). Baseline Reference: support vector machines (SVM) (cristianini2002support); random forest (RF) (breiman1996bagging); gradient boosting machine (GBM) (friedman2001greedy); k-nearest neighbor (KNN) (fix1985discriminatory); message passing neural network (MPNN) (gilmer2017neural); graph attention network (GAT) (velivckovic2017graph); graph convolutional network (GCN) (kipf2016semi); AFP (xiong2019pushing); Convolutional Neural NetworkLong Short-Term Memory (CNN, specifically Maxsmi)  (kimber2021maxsmi); Long Short-Term Memory (LSTM, specifically CLM)  (moret2022perplexity); Transformer (specifically Chemberta) (chithrananda2020chemberta); KPGT (li2023knowledge)
MoleculeACE AVG-RMSE (Ranks 1–24) MoleculeACE AVG-RMSE (Ranks 25–48)
Method AVG-RMSE $\downarrow$ Rank Method AVG-RMSE $\downarrow$ Rank
CamS-LLaMA 0.624 1 EdgePredSup 0.764 25
KPGT 0.633 2 ContextpredSup 0.764 26
CamS-LLaMA (Vocab-1K Only) 0.641 3 GraphMVP 0.768 27
CamS-LLaMA (Vocab-67K Only) 0.649 4 Infomax 0.774 28
CamS-LLaMA (w/o FP) 0.650 5 Edgepred 0.774 29
SVMECFP 0.675 6 ImageMol 0.797 30
GROVER 0.680 7 3DInfomax 0.804 31
GBMECFP 0.701 8 GraphLoG 0.807 32
RFECFP 0.705 9 KNNMACCS 0.818 33
MolFormer 0.740 10 GEM 0.821 34
KNNECFP 0.741 11 Transformer 0.868 35
MLPECFP 0.742 12 RFPHYSCHEM 0.890 36
GBMMACCS 0.742 13 GBMPHYSCHEM 0.901 37
LSTM 0.744 14 KNNPHYSCHEM 0.923 38
Contextpred 0.747 15 SVMPHYSCHEM 0.935 39
RFMACCS 0.753 16 CNN 0.937 40
GraphMAE 0.753 17 MPNN 0.959 41
SVMMACCS 0.754 18 AFP 0.970 42
MaskingSup 0.755 19 RFWHIM 0.977 43
MoleBERT 0.755 20 GBMWHIM 0.992 44
InfomaxSup 0.756 21 SVMWHIM 1.003 45
JOAO 0.757 22 GCN 1.010 46
Masking 0.758 23 KNNWHIM 1.020 47
GraphCL 0.760 24 GAT 1.049 48
Appendix E Interpretability Details for Activity-Cliff Attention Analysis
Overview. This section corresponds to the activity-cliff attention analysis in Sec. 4.3. It supplements the main text with implementation-level details of (1) activity-cliff pair construction and differential/shared atom identification (Appendix E.1), Alg.7, (2) mapping atom-level diff/shared labels to CamS tokens within each scale region (Appendix E.2), and (3) attention extraction and metric computation/aggregation for MDTA/MSTA and Rel-DTAP (Appendix E.3).
E.1 Activity-cliff Pair Construction and Differential/Shared Atom Identification
Pair construction (MoleculeACE-style).
For each MoleculeACE sub-task dataset, we start from its per-molecule CSV containing smiles, exp_mean [nM], cliff_mol, and the split label. We treat molecules with cliff_mol=1 as anchors (by default restricted to the test split), and search partners among molecules with cliff_mol=0 (by default from the same split). A candidate pair is kept if it satisfies both (1) high structural similarity and (2) large potency change. Structural similarity is a “soft-consensus” rule: we compute (a) full-molecule ECFP Tanimoto, (b) generic Murcko-scaffold ECFP Tanimoto, and (c) SMILES Levenshtein similarity, and accept the pair if any of them is $\geq\tau_{\mathrm{sim}}$ (default $0.9$). Potency change is measured by the fold change on linear exp_mean [nM] values: $\mathrm{FC}=\frac{\max(y_{a},y_{p})}{\max(\min(y_{a},y_{p}),\epsilon)}$ (with $\epsilon{=}10^{-12}$), and we keep the pair if $\mathrm{FC}\geq\tau_{\mathrm{fold}}$ (default $10$). Each selected pair yields one record $(\text{anchor},\text{partner})$; for downstream attention statistics we count the two molecules (anchor and partner) separately, as stated in Sec. 4.3.
Differential vs. shared atoms.
Given an (anchor, partner) pair, we identify shared atoms via an RDKit maximum common substructure (MCS) query with ring constraints and chemistry-aware matching (elements compared by type; bonds compared by order; completeRingsOnly and ringMatchesRingOnly enabled). Let $\mathcal{A}$ and $\mathcal{B}$ be the atom-index sets of the two molecules and let $\mathcal{M}\subseteq\mathcal{A}$ and $\mathcal{M}^{\prime}\subseteq\mathcal{B}$ be the matched atom indices returned by the first substructure match of the MCS query. We define differential atoms as the complement sets $\mathcal{A}_{\Delta}=\mathcal{A}\setminus\mathcal{M}$ and $\mathcal{B}_{\Delta}=\mathcal{B}\setminus\mathcal{M}^{\prime}$, and shared atoms as $\mathcal{M}$ and $\mathcal{M}^{\prime}$. If MCS fails (or a SMILES cannot be parsed), we conservatively treat all atoms as differential.
Atom correspondence (optional for visualization).
For visualization/debugging we also record the ordered atom correspondence induced by the MCS query, i.e., a list of paired indices $\{(i,j)\}$ obtained by aligning the two match tuples in the MCS SMARTS query order. This mapping is not required for computing Rel-DTAP, but enables atom-level highlighting across the two molecules.
E.2 Mapping Atom-level Diff/Shared Labels to Tokens in Each Scale Region
Token-to-atom alignment.
We re-encode each molecule with MultiGraphBPETokenizerExplain, which returns, for every token position $t$, an atom-index set $S_{t}$ indicating which RDKit atom indices are covered by that token (special tokens such as [BOS], [EOS], and [CONCAT] have empty sets). Because GraphBPE encoding inserts auxiliary connector atoms (*) with indices outside the original RDKit atom range, we remove such indices by clipping $S_{t}$ to $\{0,\dots,N{-}1\}$ where $N$ is the number of atoms in the original molecule.
From atom-level labels to token-level labels.
Let $\mathcal{A}_{\Delta}$ be the differential-atom set for this molecule from Appendix E.1. We assign a token-level differential indicator by an any-diff rule: a token at position $t$ is marked as differential iff $S_{t}\cap\mathcal{A}_{\Delta}\neq\varnothing$; otherwise it is marked as shared/non-differential. Thus a token covering multiple atoms is counted as differential if it contains at least one differential atom.
Scale-region boundaries in the concatenated sequence.
For multi-scale CamS sequences, scale regions are delimited by the [CONCAT] separators. Concretely, given the full token-id list (including [BOS], [EOS], and [CONCAT]), we split it into consecutive spans between [CONCAT] tokens, and exclude [BOS]/[EOS]/[CONCAT] themselves from the per-region statistics. This yields the four scale regions (e.g., 1K/7K/27K/67K in the paper) used to report scale-wise attention preferences.
E.3 Attention Extraction and Metric Computation
Attention extraction protocol.
For each molecule, we run a forward pass with output_attentions=True and extract the last-layer attention tensor. We average over heads to obtain a single attention matrix $\bar{\mathbf{A}}\in\mathbb{R}^{S\times S}$, where $S$ is the sequence length. We use the attention distribution of the final token (the last position in the sequence) as a saliency-like weighting over the prefix tokens, i.e., the row vector $\bar{\mathbf{A}}_{S,:}$. We compute this distribution in two modes: (1) without fingerprint input by running the model without the prepended fingerprint token; (2) with fingerprint input by prepending the fingerprint embedding (sequence length becomes S+1S{+}1), extracting $\bar{\mathbf{A}}_{S{+}1,:}$, and then taking only the sub-vector over molecular tokens (excluding the fingerprint position) for diff/shared statistics.
MDTA/MSTA within a scale region.
Within a given scale region $s$ (defined by [CONCAT] boundaries), let $p_{i}$ denote the (renormalized) attention weight on token $i$ in that region, and let $d_{i}\in\{0,1\}$ be its differential indicator from Appendix E.2. We define the mean differential-token attention and mean shared-token attention as
$\mathrm{MDTA}_{s}=\frac{\sum_{i}p_{i}d_{i}}{\sum_{i}d_{i}},\qquad\mathrm{MSTA}_{s}=\frac{\sum_{i}p_{i}(1-d_{i})}{\sum_{i}(1-d_{i})},$
with the convention that the corresponding mean is set to 0 if the denominator is 0 (no tokens of that type in the region).
Rel-DTAP computation and aggregation.
For each molecule (anchor and partner counted separately), we compute
$\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100,$
using $\epsilon=10^{-12}$ for numerical stability. We report the final Rel-DTAP by averaging this quantity over all molecules from activity-cliff pairs, both for each scale region $s$ and for the full concatenated sequence (all regions combined).
Algorithm 7 Shared/differential fragment labeling for an activity-cliff pair
1:Input: SMILES pair $(s_{a},s_{b})$; RDKit MCS; tokenizer-explain encoder $\iota$ returning per-token atom sets
2:Output: shared/diff atom sets $(M_{a},\Delta_{a})$ and $(M_{b},\Delta_{b})$; token-level diff masks $\mathbf{d}^{(a)},\mathbf{d}^{(b)}$; optional atom map $\mathcal{P}$
[…]
5:     completeRingsOnly, ringMatchesRingOnly,
5:     CompareElements, CompareOrder).smartsString
6:if smartssmarts is empty then
7:  $M_{a}\leftarrow\varnothing$; $M_{b}\leftarrow\varnothing$
8:else
9:  $q\leftarrow\mathrm{MolFromSmarts}(smarts)$
10:  $\mathbf{t}_{a}\leftarrow m_{a}.\mathrm{GetSubstructMatch}(q)$; $\mathbf{t}_{b}\leftarrow m_{b}.\mathrm{GetSubstructMatch}(q)$ (take the first match)
11:  $M_{a}\leftarrow\mathrm{set}(\mathbf{t}_{a})$; $M_{b}\leftarrow\mathrm{set}(\mathbf{t}_{b})$
12:  $\mathcal{P}\leftarrow\{(\mathbf{t}_{a}[i],\mathbf{t}_{b}[i])\}_{i=1}^{|\mathbf{t}_{a}|}$ (ordered atom correspondence)
13:end if
[…]
15:for side $m\in\{a,b\}$ do
16:  Encode $s_{m}$ with $\iota$ to get tokens $(x^{(m)}_{1},\dots,x^{(m)}_{L_{m}})$ and atom sets $\{S^{(m)}_{t}\}_{t=1}^{L_{m}}$
17:  Clip connector atoms: $S^{(m)}_{t}\leftarrow S^{(m)}_{t}\cap\{0,\dots,N_{m}{-}1\}$ for all $t$
18:  for $t=1$ to $L_{m}$ do
19:   if $x^{(m)}_{t}$ is [BOS] or [EOS] or [CONCAT] then
20:    $d^{(m)}_{t}\leftarrow 0$
21:   else
22:    if $S^{(m)}_{t}\cap\Delta_{m}\neq\varnothing$ then
23:     $d^{(m)}_{t}\leftarrow 1$ (diff fragment token; any-diff rule)
24:    else
25:     $d^{(m)}_{t}\leftarrow 0$ (shared fragment token)
26:    end if
27:   end if
28:  end for
29:end for